# Optimal error analysis of an interior penalty virtual element method for fourth-order singular perturbation problems

### v1

@@ Abstract
In recent studies [13, 9], the Interior Penalty Virtual Element Method (IPVEM) has been developed for solving a fourth-order singular perturbation problem, with uniform convergence established in the lowest-order case concerning the perturbation parameter. However, the resulting uniform convergence rate is only of half-order, which is suboptimal. In this work, we demonstrate that the proposed IPVEM in fact achieves optimal and uniform error estimates, even in the presence of boundary layers. The theoretical results are substantiated through extensive numerical experiments, which confirm the validity of the error estimates and highlight the method’s effectiveness for singularly perturbed problems.
@@ 1 Introduction
This paper focuses on Virtual Element Methods (VEMs), a generalization of the standard finite element method that accommodates general polytopal meshes [2, 1, 3]. Recently, the interior penalty technique for VEMs has been explored in [14] for the biharmonic equation, using the same degrees of freedom (DoFs) as for the $H^{1}$-conforming virtual elements [2]. This numerical scheme can be viewed as a combination of the virtual element space and a discontinuous Galerkin method, as the resulting global discrete space is not $C^{0}$-continuous and an interior penalty formulation is adopted to enforce the continuity of the solution. The interior penalty method is then applied to tackle the fourth-order singular perturbation problem in [13], with certain adaptations to the original space. These modifications include changes in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast to the approach in [13], we in [9] adopt the original formulation introduced in [14] to address the fourth-order singular perturbation problem (1.1). To streamline the implementation, we modify the jumps and averages in the penalty term by incorporating the elliptic projector $\Pi_{h}^{\nabla}$.
@@
Both analyses in [13, 9] demonstrate that the IPVEM is robust with respect to the perturbation parameter in the lowest-order case. However, the uniform convergence rate is only $\mathcal{O}(h^{1/2})$, which is suboptimal. To achieve optimal convergence, we can apply the widely used Nitsche’s technique, a method commonly employed for parameter-dependent boundary value problems to enhance performance. Nitsche’s technique weakly enforces boundary conditions, imposing fewer restrictions on the boundary of the domain [8]. Upon reviewing this approach, we find that the IPVEM can be viewed as an application of Nitsche’s technique, which should yield the optimal convergence rate. Building on this insight, we show in this article that the IPVEM in [9] naturally achieves optimal convergence for fourth-order singularly perturbed problems with boundary layers.
@@ 2 The continuous variational problem
We first introduce some notations. For a bounded Lipschitz domain $D$ of dimension d(d=1,2)d~(d=1,2), the symbol $(v,w)_{D}=\int_{D}vw\mathrm{d}x$ denotes the $L^{2}$-inner product on $D$, $\|\cdot\|_{0,D}$ or $\|\cdot\|_{D}$ denotes the $L^{2}$-norm. If $D=\Omega$, we abbreviate $\|\cdot\|_{\Omega}$ as $\|\cdot\|$. For a vectorial functions $\bm{v}=(v_{1},v_{2})^{T}$ and $\bm{w}=(w_{1},w_{2})^{T}$, the inner product is defined as $(\bm{v},\bm{w})_{D}=\int_{D}(v_{1}w_{1}+v_{2}w_{2})\mathrm{d}x$. The set of scaled monomials $\mathbb{M}_{r}(D)$ is given by $\mathbb{M}_{r}(D):=\{(\frac{\boldsymbol{x}-\boldsymbol{x}_{D}}{h_{D}})^{\boldsymbol{s}},~~|\boldsymbol{s}|\leq r\}$, with the generic monomial denoted by $m_{D}$, where $h_{D}$ is the diameter of $D$, $\bm{x}_{D}$ is the centroid of $D$, and $r$ is a non-negative integer. For the multi-index ${\boldsymbol{s}}\in{\mathbb{N}^{d}}$, we follow the usual notation $\boldsymbol{x}^{\boldsymbol{s}}=x_{1}^{s_{1}}\cdots x_{d}^{s_{d}}$ with $|\boldsymbol{s}|=s_{1}+\cdots+s_{d}$. Conventionally, $\mathbb{M}_{r}(D)=\{0\}$ for $r\leq-1$. Let $\mathcal{T}_{h}$ denote a tessellation of $\Omega$ into polygons. For a generic element $K$, define $h=\max_{K\in\mathcal{T}_{h}}h_{K}$ and $h_{K}={\rm diam}(K)$. The set $\mathcal{E}_{h}$ represents all the edges in $\mathcal{T}_{h}$. Let $e\subset\partial K$ be the common edge for elements $K=K_{-}$ and $K_{+}$, and let $v$ be a scalar function defined on $e$. We introduce the jump and average of $v$ on $e$ by $v]=v^{-}-v^{+}$ and $\{v\}=\frac{1}{2}(v^{-}+v^{+})$, where $v^{-}$ and $v^{+}$ are the traces of $v$ on $e$ from the interior and exterior of $K$, respectively. On a boundary edge, $[v]=v$ and $\{v\}=v$. For a non-negative integer $m$ and an element $K\in\mathcal{T}_{h}$, denote by $\mathbb{P}_{m}(K)$ the set of all polynomials on $K$ with the total degree no more than $m$. The broken Sobolev space is denoted as
@@
Moreover, for any two quantities $a$ and $b$, $a\lesssim b$” indicates $a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and $a\eqsim b$” abbreviates $a\lesssim b\lesssim a$”. The variational formulation of (1.1) is to find $u\in V:=H_{0}^{2}(\Omega)$ such that
@@
Based on this usual geometric assumptions, by the standard Dupont-Scott theory [5], for all $v\in H^{l}(K)$, there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
@@ 3 The interior penalty virtual element method
For the equipped degree of freedoms (DoFs), we refer the reader to [14, 9]. Given $v_{h}\in\widetilde{V}_{k+2}(K)$, Ref. [14] proposes an approximate $H^{1}$-elliptic projection $\Pi_{K}^{\nabla}v_{h}\in\mathbb{P}_{k}(K)$ described by the following equations:
@@
for any given $v\in V_{k}(K)$, where the quasi-average $\widehat{v}$ is defined by
@@
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big\}\Big[\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big]\mathrm{d}s,$
@@ 4 The optimal and uniform error estimate
Combining above inequalities, we derive
@@
The remaining part is to estimate the consistency term $E_{h}(u)$. By Lemma 4.1, we need to consider the estimates of $E_{A1}(u,v_{h})$, $E_{A2}(u,v_{h})$, $E_{A3}(u,v_{h})$, $\varepsilon^{2}J_{1}(I_{h}u,v_{h})$ and $\varepsilon^{2}J_{3}(I_{h}u,v_{h})$.
@@
Suppose that $\Omega$ is a bounded convex polygonal domain. Under the condition of Lemma 4.1, there holds $E_{h}(u)\lesssim\varepsilon^{1/2}\|f\|$.
@@
$\lesssim\sum_{e\in\mathcal{E}_{h}}\|w-\Pi_{0,K}^{k-2}w\|_{e}\Big\|\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\Big\|_{e}$
@@
$\lesssim\varepsilon|u|_{2}|v_{h}|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|\|v_{h}\|_{\varepsilon,h}.$
@@
$\varepsilon^{2}J_{3}(I_{h}u,v_{h})$ $=\varepsilon^{2}\sum_{e\in\mathcal{T}_{h}}\int_{e}\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\mathrm{d}s$
$\lesssim\varepsilon^{2}\sum_{e\in\mathcal{T}_{h}}\Big\|\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big\|_{0,e}\Big\|\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\Big\|_{0,e}$
@@ 5 Numerical examples
In this section, we report the performance with several examples by testing the accuracy and the robustness with respect to the singular parameter $\varepsilon$ to resolve solutions with boundary layer effects. Unless otherwise specified, we only consider the lowest-order element ($k=2$) and the domain $\Omega$ is taken as the unit square $(0,1)^{2}$.
@@
The solution with a layer is constructed such that $u(x,y)=\varepsilon(\exp(-x/\varepsilon)+\exp(-y/\varepsilon))-x^{2}y$. The right-hand side $f(x,y)=2y$ is independent of the singular parameter $\varepsilon$.
@@
To evaluate the accuracy of the proposed method, we analyze a series of meshes based on a Centroidal Voronoi Tessellation of the unit square, with polygonal counts of 32, 64, 128, 256, and 512. These meshes are generated using the MATLAB toolbox PolyMesher [12]. The convergence orders of the errors with respect to the mesh size $h$ are presented in Table 1. It is clear that for small $\varepsilon$, a quadratic convergence rate is observed, which is consistent with the theoretical prediction in Theorem 4.1.
Tab. 1: The convergence rate for Example 5.1
@@
1e-6 6.4009e-03 3.2484e-03 1.3883e-03 7.1140e-04 3.6204e-04 2.10
1e-8 6.4009e-03 3.2484e-03 1.3883e-03 7.1140e-04 3.6204e-04 2.10
@@
In this example, the exact solution takes on a more intricate form, defined as
@@
We evaluate the performance on an L-shaped domain (refer to Fig. 1 for a snapshot of the numerical and exact solutions) with $k=3$. The meshes are composed of 100, 200, 300, 400, and 500 polygons, with the corresponding results presented in Table 2. We observe a third-order convergence rate for small $\varepsilon$, consistent with the theoretical prediction in Theorem 4.1.
Tab. 2: The convergence rate for Example 5.2 on an L-shaped domain
@@
1e-6 1.1531e-04 3.8054e-05 2.0368e-05 1.3161e-05 9.7283e-06 3.13
1e-8 1.1531e-04 3.8054e-05 2.0368e-05 1.3161e-05 9.7283e-06 3.13
Fig. 1: Numerical and exact solutions for Example 5.2 ($\varepsilon=10^{-10}$)
Example 5.3.
For the Poisson problem (4.6), the exact solution is taken as $u^{0}(x,y)=\sin(\pi x)\sin(\pi y)$. Then we compute the right-hand side of (1.1) by $f=-\Delta u^{0}=2\pi^{2}\sin(\pi x)\sin(\pi y)$. The solution $u$ of the singular perturbation problem displays strong boundary layers for very small $\varepsilon$, but the solution $u$ is unknown. So we focus on the error [11]: $\|u^{0}-u_{h}\|_{\varepsilon,h}\lesssim(\varepsilon^{1/2}+h^{2})\|f\|$.
In this example, we repeat the numerical simulation using distorted meshes (cf. [9]). For very small $\varepsilon$, we can ensure a second-order convergence for the error $\|u^{0}-u_{h}\|_{\varepsilon,h}$. By assessing the relative error in (5.1) with $u$ replaced by $u^{0}$, we consistently observe second-order convergence in this example (see Table 3).
Tab. 3: The convergence rate for Example 5.3
$\N\varepsilon\backslash$ N 32 64 128 256 512 Rate
1e-8 8.6803e-03 4.3553e-03 2.2699e-03 1.1487e-03 5.7240e-04 1.95
1e-10 8.6803e-03 4.3553e-03 2.2699e-03 1.1487e-03 5.7240e-04 1.95

### v2

@@ Abstract
In recent studies [21, 15], the Interior Penalty Virtual Element Method (IPVEM) has been developed for solving a fourth-order singular perturbation problem, with uniform convergence established in the lowest-order case concerning the perturbation parameter. However, the resulting uniform convergence rate is only of half-order, which is suboptimal. In this work, we demonstrate that the proposed IPVEM in fact achieves optimal and uniform error estimates, even in the presence of boundary layers. The theoretical results are substantiated through serveral numerical experiments, which confirm the validity of the error estimates and highlight the method’s effectiveness for singularly perturbed problems.
@@ 1 Introduction
This paper focuses on the Virtual Element Method (VEM), which extends the standard finite element method to accommodate general polytopal meshes [2, 1, 3]. Compared to traditional finite element methods, VEMs offer advantages in handling complex geometries and problems with high-regularity solutions [7, 22, 25, 10]. Moreover, thanks to the great flexibility in mesh choice, VEMs require no special treatment for hanging nodes, which provides a significant advantage in mesh refinement [12, 8]. Recently, the interior penalty technique for VEMs has been explored in [24] for the biharmonic equation, using the same degrees of freedom (DoFs) as for the $H^{1}$-conforming virtual elements [2]. This numerical scheme can be viewed as a combination of the virtual element space and a discontinuous Galerkin method, as the resulting global discrete space is $H^{1}$-nonconforming and an interior penalty formulation is adopted to enforce the continuity of the solution. The interior penalty method is then applied to tackle the fourth-order singular perturbation problem in [21], with certain adaptations to the original space. These modifications include changes in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast to the approach in [21], we in [15] adopt the original formulation introduced in [24] to address the fourth-order singular perturbation problem (1.1). To streamline the implementation, we modify the jumps and averages in the penalty term by incorporating the elliptic projector $\Pi_{h}^{\nabla}$. It is worth noting that although the IPVEM is $H^{1}$-nonconforming and thus requires lower regularity assumptions on the solution [23] compared to $C^{0}$-continuous finite elements, the a priori error estimates in [24, 15, 21] show that the interior penalty virtual elements behave similarly to $C^{0}$-continuous elements. This property can be extended to the a posteriori setting, as neither the discrete scheme nor the error estimators require explicit accounting of function jumps [14].
Both analyses in [21, 15] demonstrate that the IPVEM is robust with respect to the perturbation parameter in the lowest-order case. Although two methods achieve same uniform convergence rate of $\mathcal{O}(h^{1/2})$ in theory, their numerical performance differs due to the different handling of the consistency term $\sum_{e\in\mathcal{E}_{h}}\int_{e}\frac{\partial u}{\partial\boldsymbol{n}_{e}}[v_{h}]\mathrm{d}s$. As noted in the introduction of [15] and the remark below Theorem 4.5 of [21], the nonconformity error is not optimal for the method in [21], leading to a suboptimal convergence rate as $\varepsilon$ tends to zero. This is observed in Table 4 of that paper, where a convergence rate of $\mathcal{O}(h^{1.46})$ is reported, which is better than the theoretical prediction $\mathcal{O}(h^{1/2})$ but still falls short of the optimal rate $\mathcal{O}(h^{2})$ for small $\varepsilon$. In contrast, thanks to the vanishing consistency term, our method proposed in [15] – despite also having a suboptimal error analysis – already exhibits optimal rates in numerical experiments. The present paper essentially closes the gap between theory and previously observed numerical results.
Among the discrete methods that achieve optimal and uniform convergence for singularly perturbed problems, Nitsche’s technique is commonly employed for parameter-dependent boundary value problems to enhance performance. For an overview of its applications, we refer to the review by [11]. In [17, 18], the authors show that the weak Galerkin finite element and discontinuous Galerkin methods can achieve optimal convergence by employing specially designed meshes. Since Nitsche’s method requires additional stabilization terms, which may complicate implementation, [16] reformulates the problem as a second-order system, which allows the boundary condition $\partial_{\bm{n}}u$ to be naturally and weakly imposed in the distributional mixed formulation. Meanwhile, [13] decouples the fourth-order elliptic singular perturbation problem into a combination of two Poisson equations and a generalized singularly perturbed Stokes-type equation involving the curl operator. To address the boundary layer, they introduce the nodal interpolation operator associated with the lowest-order Nédélec element of the second kind into the discrete bilinear forms, which similarly avoids additional stabilization terms.
@@
Upon reviewing this approach, we find that the IPVEM can be interpreted as an application of Nitsche’s technique. In Nitsche’s method, the boundary condition $\partial_{\bm{n}}u$ is imposed weakly via a penalty term, thereby imposing fewer restrictions on the boundary of the domain [11]. The IPVEM follows the same line, which becomes more apparent when examining the derivation of the scheme for inhomogeneous boundary conditions in [14]. Consequently, the IPVEM should yield the optimal convergence rate, as previously observed in numerical tests. Building on this insight, we demonstrate in this article that the IPVEM developed in [15] naturally achieves optimal convergence for fourth-order singularly perturbed problems with boundary layers.
@@ 2 The continuous variational problem
We first introduce some notations. For a bounded Lipschitz domain $D$ of dimension d(d=1,2)d~(d=1,2), the symbol $(v,w)_{D}=\int_{D}vw\mathrm{d}x$ denotes the $L^{2}$-inner product on $D$, $\|\cdot\|_{0,D}$ or $\|\cdot\|_{D}$ denotes the $L^{2}$-norm. If $D=\Omega$, we abbreviate $\|\cdot\|_{\Omega}$ as $\|\cdot\|$. For vectorial functions, $\bm{v}=(v_{1},v_{2})^{T}$ and $\bm{w}=(w_{1},w_{2})^{T}$, the inner product is defined as $(\bm{v},\bm{w})_{D}=\int_{D}(v_{1}w_{1}+v_{2}w_{2})\mathrm{d}x$. The set of scaled monomials $\mathbb{M}_{r}(D)$ is given by $\mathbb{M}_{r}(D):=\{(\frac{\boldsymbol{x}-\boldsymbol{x}_{D}}{h_{D}})^{\boldsymbol{s}},~~|\boldsymbol{s}|\leq r\}$, with the generic monomial denoted by $m_{D}$, where $h_{D}$ is the diameter of $D$, $\bm{x}_{D}$ is the centroid of $D$, and $r$ is a non-negative integer. For the multi-index ${\boldsymbol{s}}\in{\mathbb{N}^{d}}$, we follow the usual notation $\boldsymbol{x}^{\boldsymbol{s}}=x_{1}^{s_{1}}\cdots x_{d}^{s_{d}}$ with $|\boldsymbol{s}|=s_{1}+\cdots+s_{d}$. Conventionally, $\mathbb{M}_{r}(D)=\{0\}$ for $r\leq-1$. Let $\mathcal{T}_{h}$ denote a tessellation of $\Omega$ into polygons. For a generic element $K$, define $h=\max_{K\in\mathcal{T}_{h}}h_{K}$ and $h_{K}={\rm diam}(K)$. The set $\mathcal{E}_{h}$ represents all the edges in $\mathcal{T}_{h}$. Let $e\subset\partial K$ be the common edge for elements $K=K_{-}$ and $K_{+}$, and let $v$ be a scalar function defined on $e$. We introduce the jump and average of $v$ on $e$ by $v]=v^{-}-v^{+}$ and $\{v\}=\frac{1}{2}(v^{-}+v^{+})$, where $v^{-}$ and $v^{+}$ are the traces of $v$ on $e$ from the interior and exterior of $K$, respectively. On a boundary edge, $[v]=v$ and $\{v\}=v$. For a non-negative integer $m$ and an element $K\in\mathcal{T}_{h}$, denote by $\mathbb{P}_{m}(K)$ the set of all polynomials on $K$ with the total degree no more than $m$. The broken Sobolev space is denoted as
@@
Moreover, for any two quantities $a$ and $b$, $a\lesssim b$” indicates $a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$ and $\varepsilon$, and $a\eqsim b$” abbreviates $a\lesssim b\lesssim a$”. The variational formulation of (1.1) is to find $u\in V:=H_{0}^{2}(\Omega)$ such that
@@
Based on this usual geometric assumption, by the standard Dupont-Scott theory [4], for all $v\in H^{l}(K)$, there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
@@ 3 The interior penalty virtual element method
For the equipped degree of freedoms (DoFs), we refer the reader to [24, 15]:
  • 1.
    $\tilde{\bm{\chi}}^{p}:$ the values of $v$ at the vertices of $K$,
    $\tilde{\chi}_{z}^{p}(v)=v(z),\quad\mbox{$z$ is a vertex of $K$}.$
  • 2.
    $\tilde{\bm{\chi}}^{g}$ : the values of $h_{z}\nabla v$ at the vertices of $K$,
    $\tilde{\chi}_{z}^{g}(v)=h_{z}\nabla v(z),\quad\mbox{$z$ is a vertex of $K$},$
    where $h_{z}$ is a characteristic length attached to each vertex $z$.
  • 3.
    $\tilde{\bm{\chi}}^{e}$ : the moments of $v$ on edges up to degree k2k-2,
    $\tilde{\chi}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-2}(e),\quad e\subset\partial K.$
  • 4.
    $\tilde{\bm{\chi}}^{n}$ : the moments of $\partial_{\bm{n}_{e}}v$ on edges up to degree k1k-1,
    ${\tilde{\chi}_{e}^{n}(v)}=(m_{e},\partial_{\bm{n}_{e}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-1}(e),\quad e\subset\partial K.$
  • 5.
    $\tilde{\bm{\chi}}^{K}$ : the moments on element $K$ up to degree $k$,
    $\tilde{\chi}_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k}(K).$
Given $v_{h}\in\widetilde{V}_{k+2}(K)$, Ref. [24] proposes an approximate $H^{1}$-elliptic projection $\Pi_{K}^{\nabla}v_{h}\in\mathbb{P}_{k}(K)$ described by the following equations:
@@
for any given $v\in V_{k}(K)$, where $a^{K}(\cdot,\cdot)=(\nabla^{2}\cdot,\nabla^{2}\cdot)_{K}$ and the quasi-average $\widehat{v}$ is defined by
@@
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big\}\Big[\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big]\mathrm{d}s.$
@@ 4 The optimal and uniform error estimate
We first examine the interpolation and projection errors. In our previous work [15], we followed the proof in [22, Theorem 4.5], where the approximation errors are expressed in terms of the mesh size $h$ and yield $\mathcal{O}(h^{1/2})$ under the $H^{2}$-regularity of $u^{0}$. However, by measuring the errors with respect to $\varepsilon$ (which is relevant for the small $\varepsilon$ regime) and assuming higher regularity of $u^{0}$, we obtain optimal approximation errors.
@@
As observed, we measure the approximation errors in terms of $\varepsilon$ and assume higher regularity of $u^{0}$. Combining the above inequalities, we derive
@@
The remaining part is to estimate the consistency term $E_{h}(u)$. By Lemma 4.1, we need to consider the estimates of $E_{A1}(u,v_{h})$, $E_{A2}(u,v_{h})$, $E_{A3}(u,v_{h})$, $\varepsilon^{2}J_{1}(I_{h}u,v_{h})$ and $\varepsilon^{2}J_{3}(I_{h}u,v_{h})$. Following the same spirit as for the approximation errors, we examine the consistency error in terms of the perturbation parameter $\varepsilon$, which yields its optimal convergence when $\varepsilon\lesssim h$.
@@
Suppose that $\Omega$ is a bounded convex polygonal domain. Under the condition of Lemma 4.1, there holds $E_{h}(u)\lesssim\varepsilon^{1/2}\|f\|+{h^{k}|f|_{k-1}|v_{h}|_{1,h}}$.
@@
$\lesssim{\varepsilon^{2}}\sum_{e\in\mathcal{E}_{h}}\|w-\Pi_{0,K}^{k-2}w\|_{e}\Big\|\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\Big\|_{e}$
@@
$\varepsilon^{2}J_{3}(I_{h}u,v_{h})$ $=\varepsilon^{2}\sum_{{e\in\mathcal{E}_{h}}}\int_{e}\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\mathrm{d}s$
$\lesssim\varepsilon^{2}\sum_{{e\in\mathcal{E}_{h}}}\Big\|\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big\|_{0,e}\Big\|\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\Big\|_{0,e}$
@@ 5 Numerical examples
In this section, we report the performance with several examples by testing the accuracy and the robustness with respect to the singular parameter $\varepsilon$ to resolve solutions with boundary layer effects. We have considered both the case of $k=2$ and the case of $k=3$ and the domain $\Omega$ is taken as the unit square $(0,1)^{2}$.
@@
Let $\Omega=[0,1]\times[0,1]$. The solution with a layer is constructed such that $u(x,y)=\sin^{2}(\pi xy)+3x^{2}y^{2}$.
@@
To evaluate the accuracy of the proposed method, we analyze a series of meshes based on a Centroidal Voronoi Tessellation of the unit square, with polygonal counts of NN= 32, 64, 128, 256, and 512. These meshes are generated using the MATLAB toolbox PolyMesher [20]. The convergence orders of the errors with respect to the mesh size $h$ for $k=2$ and $k=3$ are presented in Table 1 and Table 2, respectively. It is evident that the VEM guarantees at least $k$-th order convergence for all $\varepsilon\in(0,1]$. As $\varepsilon$ approaches zero, a $(k+1)$-th order convergence rate is observed, which agrees with the theoretical prediction in Theorem 4.1. Furthermore, the error behavior remains stable in this regime.
Tab. 1: The convergence rate with $k=2$ for Example 5.1
@@
1e-0 1.6926e-02 1.1886e-02 7.7060e-03 6.0383e-03 3.8926e-03 1.04
1e-1 6.0237e-02 3.8000e-02 2.4672e-02 1.9601e-02 1.3106e-02 1.07
1e-2 1.4605e-02 8.4552e-03 3.7313e-03 2.3571e-03 1.4715e-03 1.69
1e-3 1.2650e-02 6.8511e-03 2.7067e-03 1.3758e-03 6.9842e-04 2.13
1e-4 1.2636e-02 6.8360e-03 2.6888e-03 1.3629e-03 6.8605e-04 2.15
1e-5 1.2636e-02 6.8359e-03 2.6888e-03 1.3628e-03 6.8597e-04 2.15
Tab. 2: The convergence rate with $k=3$ for Example 5.1
$\N\varepsilon\backslash$ N 32 64 128 256 512 Rate
1e-0 3.8262e-03 1.4450e-03 6.2458e-04 2.6427e-04 1.4084e-04 2.40
1e-1 1.3452e-02 5.0626e-03 2.1361e-03 9.2470e-04 4.9631e-04 2.39
1e-2 3.0331e-03 1.2243e-03 3.7252e-04 1.5741e-04 7.5709e-05 2.72
1e-3 1.9226e-03 7.1015e-04 1.9430e-04 7.5048e-05 2.9407e-05 3.06
1e-4 1.9592e-03 7.1154e-04 1.8983e-04 6.8974e-05 2.5094e-05 3.19
1e-5 1.9598e-03 7.1187e-04 1.8989e-04 6.8994e-05 2.5104e-05 3.19
@@
In this example, we test the performance of the VEM to resolve a solution with boundary layer effect. The solution is construct such that
@@
For this boundary layer problem, we focus on the behavior when $\varepsilon$ is small. We evaluate the performance on a square domain. Please refer to Fig. 1 for a snapshot of the numerical and exact solutions with $\varepsilon=10^{-10}$. Additionally, we present the error and the corresponding convergence rate for $k=2$ and $k=3$ in Tables 3 and 4, respectively. For sufficiently small $\varepsilon$, the observed convergence rates agree with the theoretical prediction in Theorem 4.1.
Tab. 3: The convergence rate for Example 5.2 with $k=2$ on a square domain
@@
1e-4 7.8922e-03 4.0052e-03 2.0022e-03 1.0141e-03 4.9884e-04 1.99
1e-5 7.8921e-03 4.0052e-03 2.0022e-03 1.0140e-03 4.9877e-04 1.99
1e-6 7.8921e-03 4.0052e-03 2.0022e-03 1.0140e-03 4.9877e-04 1.99
1e-7 7.8921e-03 4.0052e-03 2.0022e-03 1.0140e-03 4.9877e-04 1.99
Tab. 4: The convergence rate for Example 5.2with $k=3$ on a square domain
$\N\varepsilon\backslash$ N 100 200 300 400 500 Rate
1e-4 1.1235e-03 3.8181e-04 1.3050e-04 4.5905e-05 1.6040e-05 3.06
1e-5 1.1238e-03 3.8195e-04 1.3056e-04 4.5928e-05 1.6044e-05 3.06
1e-6 1.1238e-03 3.8195e-04 1.3056e-04 4.5928e-05 1.6044e-05 3.06
1e-7 1.1238e-03 3.8195e-04 1.3056e-04 4.5928e-05 1.6044e-05 3.06
Fig. 1: Numerical and exact solutions for Example 5.2 with $k=2$ ($\varepsilon=10^{-10}$)